OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. 1, -4, -9, -14, . . .

OpenStudy (anonymous):

You need to find the difference in each term and use the formula for the nth term. \[n_{th} = a_1+(n-1)d\]

OpenStudy (anonymous):

Since you need the sum of the first 12 terms you need to add them all up once you find them.

OpenStudy (anonymous):

The sum of that can also be done by a formula.

OpenStudy (mathstudent55):

Most sequences we study are either arithmetic or geometric sequences. Are you familiar with those terms?

OpenStudy (anonymous):

\[S_{n} = \frac{ n(a_1+a_n) }{ 2 }\] Where, n = the number of terms you are summing. a1 = the first term. an = the last term.

OpenStudy (anonymous):

d= 5 ?

OpenStudy (anonymous):

You can use the first formula to find the last term.

OpenStudy (anonymous):

Almost your numbers are negative.

OpenStudy (anonymous):

And yes, MathStudent

OpenStudy (mathstudent55):

Ok, so you know this is a geometric sequence. To find d, subtract any term from the next one. For example, d = -9 - (-5) = -5

OpenStudy (mathstudent55):

By adding the common difference, you can generate the next several terms. Then add them up to find the sum. Of course, you can also use the formula @Johnbc gave you above.

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

You're welcome.